OpenStudy (anonymous):

How do I simplify the fifth sqrt of 1/32 ?

OpenStudy (nenadmatematika):

1/2

OpenStudy (kinggeorge):

\[\sqrt[5]{1 \over 32} = {1 \over \sqrt[5]{32}} = {1 \over 2}\]since \(2^5 = 32\)

OpenStudy (anonymous):

1 = 1*1*1*1*1 32 = 2*2*2*2*2 so 1/2

OpenStudy (anonymous):

Awesome! Thanks everyone!

OpenStudy (anonymous):

u r welcome...☺

OpenStudy (anonymous):

One more question... it was actually \[-\sqrt[5]{1-32}\] so does that make the answer - 1/2 ?

OpenStudy (kinggeorge):

If that was the actual equation, it should be very close to, but not exactly \(-{1 \over 2}\)

OpenStudy (anonymous):

Great, thank you :)